solve by any method you choose
x-y+3z=13
4x+y+2z=17
3x+2y+2z=1

Answer by richwmiller(17219)

x-y+3z=13
4x+y+2z=17
3x+2y+2z=1
1,-1,3,13
4,1,2,17
3,2,2,1
add down (-4/1) *row 1 to row 2
1,-1,3,13
0,5,-10,-35
3,2,2,1
add down (-3/1) *row 1 to row 3
1,-1,3,13
0,5,-10,-35
0,5,-7,-38
divide row 2 by 5/1
1,-1,3,13
0,1,-2,-7
0,5,-7,-38
add down (-5/1) *row 2 to row 3
1,-1,3,13
0,1,-2,-7
0,0,3,-3
divide row 3 by 3/1
1,-1,3,13
0,1,-2,-7
0,0,1,-1
We now have the value for the last variable.
We will work our way up and get the other solutions.
add up (2/1) *row 3 to row 2
1,-1,3,13
0,1,0,-9
0,0,1,-1
add up (-3/1) *row 3 to row 1
1,-1,0,16
0,1,0,-9
0,0,1,-1
add up (1/1) *row 2 to row 1
1,0,0,7
0,1,0,-9
0,0,1,-1
final
1,0,0,7
0,1,0,-9
0,0,1,-1

(7,-9,-1)